Circular No. 7597

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001ae IN IC 4229 A. Friedman and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 17.3) on unfiltered images taken on Mar. 15.5 and 16.3 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001ae is located at α = 13^{h}22^{m}26^{s}.16, δ = −2^{o}25′25′′.0 (equinox 2000.0), which is 1′′.0 east and 18′′.3 south of the nucleus of IC 4229. The new object was also confirmed on an unfiltered CCD image taken by M. Schwartz on Mar. 16.4 with the Tenagra Observatory 0.5-m automatic telescope. A KAIT image taken on Feb. 27.5 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001ab IN NGC 6130 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001ab (cf. IAUC 7595), obtained by M. Calkins on Mar. 14.52 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova. The spectrum consists of a blue continuum with superposed P-Cyg lines of hydrogen.

SUPERNOVA 2001ac IN NGC 3504 Matheson et al. add: “A spectrum of SN 2001ac (cf. IAUC 7596),

obtained as above by Calkins on Mar. 14.45 UT, shows an object with a blue continuum and narrow hydrogen emission lines (FWHM ∼ 900 km/s). The broader component of the Balmer lines often seen in type-IIn supernovae is not apparent. It is similar to the spectrum of the type-IIn SN 1997bs (Van Dyk et al. 2000, PASP 112, 1532) and SN 1999bw (IAUC 7150); this may be a superoutburst of a luminous blue variable rather than a true supernova.”

SUPERNOVA 2001ad IN NGC 6373 Matheson et al. also write: “A spectrum of SN 2001ad (cf. IAUC 7596),

obtained as above by P. Berlind on Mar. 15.49 UT, shows it to be a type-II supernova. The spectrum consists of a continuum with superposed P-Cyg lines of hydrogen. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3320 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is ∼ 12 000 km/s.”

2001 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green